Solve the system by substitution
5x+2y=-15
y=2x-12

Answer by funmath(2873)

Solve the system by substitution
5x+2y=-15 R1
y=2x-12 R2
Let y=(2x-12) in R1 and solve for x
5x+2(2x-12)=-15
5x+4x-24=-15
9x-24=-15
9x-24+24=-15+24
9x=9
9x/9=9/9
x=1
Let x=1 in R2 and solve for y
y=2(1)-12
y=2-12
y=-10
The solution to the system is (x,y)=(1,-10)
Happy Calculating!!!!